OpenStudy (starryknight):

what does x equal? -2x – 0y = -4 -6x – 2y = -12 x = 2 x = -2 x = -10 x = 4

OpenStudy (starryknight):

can someone brake it down for me to unde rstand

OpenStudy (whpalmer4):

\[-2x-0y=-4\]If we are multiplying \(y\) by \(0\), the result is \(0\):\[-2x-0=-4\]\[-2x=-4\]\[\frac{-2x}{-2} = \frac{-4}{-2}\]\[x=2\]Agreed?

OpenStudy (whpalmer4):

Solving for \(y\) when \(x =2\) \[-6x-2y=-12\]\[-6(2)-2y=-12\]\[-12-2y=-12\]Add 12 to both sides\[-12+12-2y=-12+12\]\[-2y=0\]\[y=0\] test our solution in both equations: \[-2(2)-0(0) = -4\]\[-4-0=-4\]\[-4=-4\checkmark\]\[-6(2) -2(0)=-12\]\[-12-0=-12\]\[-12=-12\checkmark\] You need to test in ALL of the equations to make sure you have a valid solution.

OpenStudy (starryknight):

omg thank you a million times my teacher tried to explain it to me and i couldnt understand :))))))))

OpenStudy (whpalmer4):

Happy to be of help! :-)